### v3

# Furthermore, can this Aschenbach-like phenomenon in static configurations be considered as an observable signature in the dynamics of general relativity, similar to the original Aschenbach effect in rotating spacetimes?

Abstract
The Aschenbach effect is widely regarded as a manifestation of two quintessential relativistic features: frame dragging and extreme spacetime curvature. Traditionally associated with rotating geometries, this non-monotonic behavior in orbital angular velocity challenges Newtonian intuition. In our previous work [11], however, we demonstrated that this velocity irregularity is not exclusive to spinning spacetimes. Specifically, we showed that the presence of a stable minimum in the gravitational potential, corresponding to a stable photon sphere, can reproduce Aschenbach-like behavior in static black holes as well. This observation suggests that, even in the absence of rotational frame dragging, curvature alone—if encoded through appropriate geometric extrema—may be sufficient to induce non-monotonic velocity profiles.
In this study, we build upon that foundation to investigate whether black hole architectures in theories of Massive Gravity can inherently support the emergence of Aschenbach-like phenomena.
Furthermore, can this Aschenbach-like phenomenon in static configurations be considered as an observable signature in the dynamics of general relativity, similar to the original Aschenbach effect in rotating spacetimes?
keywords: Aschenbach effect, photon spheres, Time-like Circular Orbits, black holes, Massive Gravity
Can Static Black Holes in Massive Gravity Serve as Candidates for Aschenbach-Like Phenomena?
Mohammad Ali S. Afshar  [^1] Email:   m.a.s.afshar@gmail.com,    Jafar Sadeghi  [^2] Email:   pouriya@ipm.ir,    Tahereh Azizi  [^3] Email:   t.azizi@umz.ac.ir,    A. S. Sefiedgar  [^4] Email:   a.sefiedgar@umz.ac.ir
Department of Physics, Faculty of Basic Sciences,
University of Mazandaran P. O. Box 47416-95447, Babolsar, Iran
School of Physics, Damghan University, P. O. Box 3671641167, Damghan, Iran
Canadian Quantum Research Center, 32 Ave, 204-3002 V1T 2L7, Vernon, BC, Canada
1 Introduction
1.1 The Aschenbach Effect: Significance and Origins
The Aschenbach effect represents a compelling manifestation of general relativity in astrophysical con texts. As a tangible prediction of Einstein’s theory in the strong-field regime, observing this phenomenon would confirm that spacetime near black holes behaves precisely as the field equations predict [1, 2, 3] . the systematic analysis of orbital stability near Kerr black holes gained momentum in the late 20th century. Key figures such as Brandon Carter, and later Marek A. Abramowicz, made substantial contributions to the mathematical formalism underpinning these orbital dynamics [4, 5, 6]. Accordingly, This effect was first identified in 2004 when Aschenbach discovered, through numerical simulations, an anomalous drop in orbital velocity for test particles approaching the event horizon of rapidly rotating Kerr black holes [7]. This non-monotonic behavior directly challenges traditional Keplerian flow models com monly employed in simplified accretion disk theories. Unlike quasi-periodic oscillations (QPOs), which typically require magneto-hydrodynamic (MHD) simulations for accurate modeling [8], the Aschenbach effect is purely kinematic and emerges directly from Kerr geometry. Rigorous inferences from Einstein’s field equations demonstrate that the phenomenon originates from the interplay between gravitational forces and rotational effects encoded in the Kerr metric.
1.2 Fundamental Physical Elements
Innermost Stable Circular Orbit (ISCO):
This represents the smallest radius at which a test particle can maintain a stable circular orbit around a black hole. Beyond this threshold, minor perturbations trigger inward plunges. The ISCO effectively delineates the inner edge of accretion disks, with matter interior to this radius undergoing rapid radial infall. In rotating spacetimes, ISCO properties depend significantly on orbital orientation relative to the black hole’s spin [9].
Frame Dragging:
The rotation of a Kerr black hole ”drags” surrounding spacetime, altering the dynamics of nearby matter. This effect, quantified by the spin parameter $a=J/M$ (where $J$ is angular momentum and $M$ is mass), is strongest near the event horizon. For maximal spin ($a/M=1$), frame dragging reaches its extremum; it vanishes completely for non-spinning Schwarzschild black holes ($a=0$) [9].
Strong-Field Gravity and Angular Velocity:
In Newtonian gravity, angular velocity $\Omega$ scales monotonically with radius as $\Omega_{\text{Newton}}(r)=\sqrt{GM/r^{3}}$. However, near rapidly rotating Kerr black holes, the angular velocity follows $\Omega_{\text{Kerr}}(r)=\sqrt{M}/(r^{3/2}+a\sqrt{M})$, revealing a complex radial evolution [10]:
1. At large radii: Keplerian-like increase in $\Omega$ as $r$ decreases ($d\Omega/dr<0$)
2. At intermediate radii: Frame-dragging dominates, causing $\Omega$ to rise super-Keplerianly
3. At a critical radius $r_{\text{peak}}$: Spacetime curvature overtakes rotational effects, causing $\Omega$ to decrease as $r$ decreases ($d\Omega/dr>0$)
This unexpected reversal—the peak and drop in orbital angular velocity near the event horizon—defines the Aschenbach effect and arises from the interplay between intense gravitational curvature and spin-induced frame dragging. While initially identified in asymptotically flat spacetimes, this phenomenon has also been rigorously investigated in asymptotically anti-de Sitter (AdS) and de Sitter (dS) geometries [7].
1.3 Beyond Rotation: Static Configurations and the Aschenbach-like Effect
A fundamental question:
Is non-monotonic angular velocity exclusive to rotating black holes, or can static black holes—devoid of intrinsic spin and frame dragging—reproduce similar behavior through alternative mechanisms?
In our previous work [11], we demonstrated that while static black holes lack rotational effects, they retain strong gravitational curvature. We proposed that the emergence of a stable gravitational potential minimum—either local or global—outside the event horizon, typically associated with a stable photon sphere, could serve as a surrogate mechanism replicating dynamical aspects of frame dragging.
Although historically, such ”minimum-maximum” potential structures have been well-documented in horizonless ultra-compact objects, with most naked singularity models exhibiting stable photon spheres [12, 13, 14]. However, when the Weak Cosmic Censorship Conjecture (WCCC) is enforced, introducing an event horizon typically hides these minima behind the horizon, resulting in purely monotonic velocity profiles. Nevertheless, certain models maintain exterior potential minima despite the presence of an event horizon, yielding non-monotonic angular velocity profiles centered around these extrema [15, 16].
1.4 Massive Gravity as a Testing Ground
Encouraged by findings in a hybrid Gauss-Bonnet-Massive Gravity black hole model [11] and similar results in another Massive Gravity formulation [16], we hypothesize that black hole geometries within Massive Gravity theories may inherently support Aschenbach-like phenomena. To test this hypothesis, we investigate three distinct black hole models in Massive Gravity, analyzing the conditions under which non-monotonic angular velocity profiles emerge.
Finally, It is crucial to emphasize that while the observational manifestations may resemble one another, the underlying physics differs substantially: - In rotating (Kerr) black holes, the Aschenbach effect stems from frame dragging coupled with strong gravity
- In static black holes, analogous behavior is induced through exterior gravitational potential minima, absent frame dragging
This fundamental distinction justifies our terminological precision: we refer to the phenomenon in static configurations as the ”Aschenbach-like effect” rather than the Aschenbach effect proper. Our investigation seeks to determine whether this Aschenbach-like phenomenon in static configurations can serve as an observable signature of general relativistic dynamics, similar to the original Aschenbach effect in rotating spacetimes.
2 Methodology
In this Article, for the analysis of photon spheres, we adopt the topological charge-based and approach, which has been thoroughly detailed in [17, 18] and our previous works [12, 13]. In the interest of brevity and to avoid redundancy, we do not revisit the foundational concepts here.
Also, in examining TCOs , we rely on the orbital relations and stability criteria formulated in studies [19, 20, 21]. Given the comprehensive nature of these references, we direct interested readers to those publications for an in-depth treatment of the methodology.
2.1 Topological Photon Sphere
To study the photonic sphere, we first need an effective potential. We use the effective potential that was constructed for the study of the topological method based on the metric coefficients of a spherically symmetric four-dimensional model. We have [17, 18]:
$\mathit{ds}^{2}=-\mathit{dt}^{2}f\!\left(r\right)+\frac{\mathit{dr}^{2}}{g\!\left(r\right)}+\left(d\theta^{2}+d\varphi^{2}\sin\!\left(\theta\right)^{2}\right)h\!\left(r\right)=\mathit{dr}^{2}g_{\mathit{rr}}-\mathit{dt}^{2}g_{\mathit{tt}}+d\theta^{2}g_{\theta\theta}+d\varphi^{2}g_{\varphi\varphi},$ (2.1)
$\begin{split}H(r,\theta)=\sqrt{\frac{-g_{tt}}{g_{\varphi\varphi}}}=\frac{1}{\sin\theta}\bigg(\frac{f(r)}{h(r)}\bigg)^{1/2}.\end{split}$ (2.2)
Now, to determine the charges and topological behavior of the model, we will first need a vector field using the effective potential $H$. For this purpose, we consider a general two dimensional vector field as $\phi$ which can be decomposed into two components, $\phi^{r}$ and $\phi^{\Theta}$ [17, 18],
$\phi=(\phi^{r},\phi^{\Theta}),$ (2.3)
[…]
where $\iota=1,2$ and $(\phi^{1}=\phi^{r})$ , $(\phi^{2}=\phi^{\Theta})$. In terms of $H$, the vector field $\phi=(\phi^{r},\phi^{\Theta})$ can be written as follows:
$\begin{split}&\phi^{r}=\frac{\partial_{r}H}{\sqrt{g_{rr}}}=\sqrt{g(r)}\partial_{r}H,\\ &\phi^{\Theta}=\frac{\partial_{\theta}H}{\sqrt{g_{\theta\theta}}}=\frac{\partial_{\theta}H}{\sqrt{h(r)}}.\end{split}$ (2.5)
By employing the previously established relations and the definition of topological charge, one can systematically assign a charge to each photon sphere present in the spacetime [12, 13, 18]. Before drawing broader conclusions from this section, it is important to underscore a key finding from our earlier publications: In configurations where the model exhibits black hole characteristics, typically satisfying the condition of possessing an event horizon, one often finds a single unstable photon sphere located outside the horizon, characterized by a net topological charge of 1-1. From the perspective of the effective potential, this corresponds to the presence of at least one maximum outside the horizon. Conversely, in horizonless configurations (i.e., naked singularities), the geometry commonly features not only an unstable photon sphere with a topological charge of 1-1, but also one or two stable photon spheres with topological charges of +1. These are identified as local minima of the potential structure and represent dynamically stable light rings. These distinctions are of fundamental importance and will be revisited in subsequent discussions, particularly when analyzing the relationship between topological photon sphere and orbital dynamics in strongly curved spacetimes.
2.2 Time-like Circular Orbits (TCOs)
The geodesic behavior in any gravitational model is profoundly shaped by the form of the effective potential, derived from the interplay of fields within the model’s action. These geodesics may either lack recognizable turning points or exhibit them, forming closed trajectories. In the case of null geodesics, such closed paths correspond to photon spheres, whereas in timelike geodesics, they manifest as TCOs. Both types of orbits may be either stable or unstable, depending on the curvature and potential landscape of the spacetime. In this study, we adopt a static, axisymmetric spacetime with an imposed $\mathbb{Z}_{2}$ symmetry under a (1+3) dimensional framework. This symmetry structure ensures that focusing the analysis on the equatorial plane does not compromise the generality of our results. Based on the metric equation—namely Eq. (2.1)—we consider the following key quantities [19]:
$A=g_{\varphi\varphi}E^{2}+g_{\mathit{tt}}L^{2},$ (2.6)
$B=-g_{\varphi\varphi}g_{\mathit{tt}},$ (2.7)
which $E$ and $L$ are the energy and the angular momentum respectively. Now the Lagrangian can be recast as:
$2\mathfrak{L}=-\frac{A}{B}=\varrho,$ (2.8)
where $\varrho=-1,0$ for time-like and null geodesics, respectively. Using the above Lagrangian, the effective potential can be rewritten as follows:
$V_{\mathit{eff}\!\left(\varrho\right)}\!\left(r\right)=\varrho+\frac{A}{B}.$ (2.9)
The concept of angular velocity and $\beta$ in terms of metric parameters [19].
$\Omega=\frac{g_{\mathit{tt}}L}{g_{\varphi\varphi}E},$ (2.10)
$\beta=-g_{tt}-2g_{t\varphi}\Omega-g_{\varphi\varphi}\Omega^{2}=-r^{2}\Omega^{2}+f(r),$ (2.11)
An important aspect to consider in the study of TCOs is the value of the $\beta$ function. In regions where $\beta$ is negative, the energy and angular momentum become imaginary, rendering these areas effectively forbidden for the presence of TCOs. Interested readers seeking a deeper understanding of these relationships, and more examples, are encouraged to refer to [19, 20, 21].
3 Black holes in Massive Gravity: Advantages and disadvantages
The concept of a massive graviton originated with Markus Fierz and Wolfgang Pauli. They formulated a linear theory (Fierz-Pauli theory) describing a massive spin-2 particle in flat spacetime [22, 23]. However, this theory suffered from the van Dam-Veltman-Zakharov (vDVZ) discontinuity [22, 24]: predictions didn’t smoothly reduce to General Relativity (GR) as the graviton mass approached zero, conflicting with precise solar system tests [22, 25]. Vainshtein proposed a resolution. He demonstrated that nonlinear interactions in massive gravity lead to strong effects (”Vainshtein screening”) near massive sources, dynamically suppressing the helicity-0 mode of the graviton and restoring agreement with General Relativity (GR) in the massless limit, despite a finite graviton mass. This mechanism resolved the van Dam-Veltman-Zakharov (vDVZ) discontinuity by showing that linearized approximations break down in strong-field regimes, such as solar system scales. However, a consistent nonlinear completion of massive gravity—free of the Boulware-Deser ghost instability—was only achieved decades later by de Rham, Gabadadze, and Tolley (dRGT theory, 2010) [26], which explicitly incorporates Vainshtein screening as a core feature. Constructing a consistent, ghost-free nonlinear massive gravity theory proved challenging for decades due to the emergence of Boulware-Deser ghosts (unphysical negative-energy states) in early formulations. The field was revolutionized in 2010 by Claudia de Rham, Gregory Gabadadze, and Andrew Tolley (dRGT massive gravity). They identified a specific class of nonlinear interaction terms—tuned potential terms constructed from invariants of the dynamical metric and the fixed reference metric—which systematically eliminate the Boulware–Deser ghost to all orders in the Hamiltonian formulation. This provided the first consistent four-dimensional (4D), Lorentz-invariant theory of massive gravity in flat spacetime, free of the instability that had plagued earlier attempts. The theory relies on a reference metric (often Minkowski) and a carefully tuned structure of potential terms to preserve constraints that eliminate the ghost. Subsequent analyses by S.F. Hassan and R.A. Rosen confirmed the absence of the Boulware-Deser ghost in dRGT through a canonical ADM formalism [26, 27].
Shortly after the dRGT formulation, intense research began to find black hole solutions:
$\clubsuit$ Schwarzschild-like Solutions: Early solutions assumed specific symmetries or ansatzes, often recovering the standard Schwarzschild metric but with the graviton mass parameters appearing as effective cosmological constants or influencing asymptotic behavior differently than in GR.
$\clubsuit$ Hairy Black Holes: It became clear that dRGT massive gravity allows for black holes with ”hair” – additional parameters beyond mass, charge, and angular momentum, forbidden in GR by no-hair theorems. These hairs are directly related to the graviton mass parameters and the structure of the potential. Spherically symmetric solutions often involve additional fields (Stückelberg fields) or manifest as modifications to the metric components.
$\clubsuit$ Bimetric Gravity: A closely related and dynamically richer framework involves two interacting metrics: one dynamical (spacetime) and one fixed reference metric (often Minkowski or de Sitter). Of course, it should be noted that this fixed reference metric is from the perspective of the massive gravity of dRGT, which seems necessary and one can also be considered it as a dynamic reference. Bimetric gravity naturally incorporates massive gravity limits and provides a more flexible setting for finding black hole solutions, including rotating ones (analogues of Kerr) which are much harder to find in pure dRGT [27, 28].
Although the formation of a black hole in the form of Massive gravity still follows the same basic principles of general relativity, but Massive gravity modifies how gravity works. The key differences in both the collapse process and the resulting black hole structure originate from the modified gravitational dynamics of Massive Gravity. This means:
1. Modified Field Equations:
Einstein’s equations are replaced by equations where the graviton mass terms act as an effective source of stress-energy:
$G_{\mu\nu}=8\pi GT_{\mu\nu}(matter)+T_{\mu\nu}(mass).$
This $T_{\mu\nu}(mass)$ depends non-trivially on the metric and the reference/Stückelberg fields.
2. Altered Dynamics:
The mass terms suppress gravitational fluctuations at large distances (Yukawa-like decay), change the propagation speed of gravity waves, and introduce new degrees of freedom. This affects the collapse process and the equilibrium configurations.
3. Vainshtein Mechanism:
The Vainshtein mechanism is a nonlinear screening effect that dynamically suppresses the extra degrees of freedom (the 5 polarizations of the massive graviton) and the deviations from General Relativity in regions of strong gravity or high matter density. Near the collapsing object and the resulting black hole, the Vainshtein mechanism is active and making the gravitational field look very similar to General Relativity close to the horizon, while deviations become significant at larger distances (asymptotics).
4. ”Hair” Formation:
The requirement to satisfy the modified field equations and boundary conditions, often involving the Stückelberg fields or the interplay with the reference metric, necessitates additional parameters ”hair” beyond mass, charge, and angular momentum to fully characterize the black hole.
In summary:
Black holes in massive gravity are a direct consequence of gravitational collapse within a modified theory where the graviton has mass. The graviton mass terms can act as an effective dark energy driving the late-time accelerated expansion of the universe without needing a cosmological constant or quintessence fields. Driven by the resolution of the ghost problem (dRGT theory), they represent a vibrant research area.
Their key features are modified asymptotics, the potential for ”hair,” and the influence of the Vainshtein mechanism near the horizon.
Their Advantages include potential explanations for dark energy/dark matter, singularity resolution  [^5] Some solutions in massive gravity exhibit regular black holes or horizons without singularities. The effective stress-energy from the mass terms can violate energy conditions, potentially halting collapse before a singularity forms. This could resolve the central singularity problem of GR, and unique observational signatures.
However, significant disadvantages persist: theoretical complexity, stability concerns, stringent constraints from gravitational wave astronomy, causality and superluminality  [^6] Some solutions exhibit apparent superluminal propagation of fluctuations around the background solutionand naturalness [^7] Explaining the extremely small graviton mass required by observations (if non-zero) is challenging from a particle physics perspective issues.
3.1 Bardeen AdS Black Holes in Massive Gravity
In the presence of massive gravity, the action coupled to nonlinear electrodynamics with negative cosmological constant is given by [29]:
$S$ $=$ $\frac{1}{2}\int d^{4}x\sqrt{-g}\Big[R-2\Lambda+m_{g}^{2}\sum_{i}^{4}c_{i}\,\mathcal{U}_{i}(g,f)-\frac{1}{4\pi}\mathcal{L}(F)\Big],$ (3.1)
[…]
${\cal{L}}(F)=\frac{1}{2sq^{2}}\left(\frac{\sqrt{2e^{2}F}}{1+\sqrt{2e^{2}F}}\right)^{\frac{5}{2}},$
with $\left[\mathcal{K}\right]=\mathcal{K}_{\mu}^{\mu}$ , $\mathcal{L}(F)$ a function of $F\equiv\frac{1}{4}F_{\mu\nu}F^{\mu\nu}$ is the Lagrangian density of Bardeen source and the parameter $s$ is related to the magnetic monopole charge $q$ and mass $M$ by relation $s=q/2M$ [30, 31].The spherically-symmetric static metric line element of such a spacetime is given by the following [31]:
$f(r)=1-\frac{2Mr^{2}}{(r^{2}+q^{2})^{\frac{3}{2}}}+\frac{r^{2}}{l^{2}}+m_{g}^{2}\left(C^{2}c_{2}+\frac{Cc_{1}r}{2}\right),$ (3.2)
where $l$ is AdS radius.
Figure 1: Metric function Fig (1a): With $l=1,q=0.5,C=1,c_{1}=-1,c_{2}=1,m_{g}=0.5$ with different M , (1b): With $l=1,M=10,C=1,c_{1}=-1,c_{2}=1,m_{g}=0.5$ with different q (1c): With $M=2,q=0.5,C=1,c_{1}=-1,c_{2}=1,m_{g}=0.5$ with different l (1d): With $M=2,l=1,q=0.5,C=1,c_{1}=-1,c_{2}=1$ with different $m_{g}$ (1e): With $M=2,l=1,q=0.5,C=1,c_{2}=1,m_{g}=0.5$ with different $c_{1}$ for Bardeen black holes in massive gravity
As stated in the past, since the behavior of black holes in a massive gravitational structure is deeply dependent on parameter values, it is first useful to consider the effect of changes in each parameter on the radius of the event horizon. The behavioral differences of the metric function, driven by parametric variations, are clearly observable in Fig. (1). As evident from panels Fig. (1) and Fig. (1), the interplay between the black hole mass and the graviton mass significantly influences the structure of the solution. When these two masses approach one another in magnitude, the system tends toward a naked singularity regime, implying a breakdown of the event horizon. A similar trend is observed with increasing electric charge Fig. (1): higher values of charge progressively push the system toward singular behavior, eliminating the possibility of horizon formation. Furthermore, increasing the curvature radius of the spacetime—alternatively interpreted as the AdS radius—directly affects the radius of the event horizon. As depicted in Fig. (1), larger values of $l$ expand the horizon radius; however, excessive growth (e.g., $l=20$) leads to the degeneration of the metric function into a single-root structure, effectively eliminating the outer horizon. The influence of the parameter $c_{1}$ appears negligible within small domains. Nevertheless, beyond a critical threshold, its contribution becomes substantial, causing the Cauchy horizon and the event horizon to approach one another Fig. (1), which signals a reduction in the inter-horizon region.
In search of the Aschenbach-like phenomenon
Given the complexity of the model and the multitude of governing parameters, it is evident that an analytical solution is infeasible. As such, parameter initialization followed by numerical analysis is essential—an approach we were already compelled to adopt in our study of the metric function. To proceed toward detecting Aschenbach-like behavior, we now consider a baseline configuration by assigning the values $M=2,l=1,q=0.5,C=1,c_{1}=-1,c_{2}=1,m_{g}=0.5$ to the relevant parameters. This initial scenario serves as a reference point to investigate the emergence of non-monotonic angular velocity features under simplified but physically representative conditions. To ensure that the model under consideration truly exhibits black hole behavior, it is instructive to begin with an analysis of its general structural features. By numerically solving the metric function, we find that the spacetime possesses two distinct horizons:
- A Cauchy horizon at $r=0.2306$,
- An event horizon at $r=1.2232$.
This double-horizon structure satisfies the Weak Cosmic Censorship Conjecture (WCCC), confirming that any singularity remains hidden behind an event horizon. Furthermore, using the relations for mass:
$M=\frac{\left(q^{2}+r^{2}\right)^{\frac{3}{2}}\left(2C^{2}l^{2}c_{2}m_{g}^{2}+C\,l^{2}rc_{1}m_{g}^{2}+2l^{2}+2r^{2}\right)}{4r^{2}l^{2}}$ (3.3)
and temperature:
$T=\frac{\left(2C\,r^{3}c_{1}m_{g}^{2}+\left(2C^{2}c_{2}m_{g}^{2}+2\right)r^{2}-C\,q^{2}rc_{1}m_{g}^{2}-4q^{2}\left(C^{2}c_{2}m_{g}^{2}+1\right)\right)l^{2}+6r^{4}}{8r\,l^{2}\left(q^{2}+r^{2}\right)\pi}$ (3.4)
we can compute the heat capacity of the model via below Equation.
$\varsigma$ $=$ $\frac{\partial M}{\partial T}\bigg|_{r_{h}}=\left(\frac{\partial M}{\partial r}\right)_{r_{h}}\left(\frac{\partial r}{\partial T}\right)_{r_{h}}.$ (3.5)
The local thermodynamic stability of a black hole is typically assessed by examining the behavior of its heat capacity:
- A positive heat capacity corresponds to thermodynamic stability.
- A negative heat capacity indicates instability , as established in the literature [32, 33, 34].
Figure 2: The heat capacity with $M=2,l=1,q=0.5,C=1,c_{1}=-1,c_{2}=1,m_{g}=0.5$ for Bardeen black holes in massive gravity
By plotting the heat capacity equation with respect to the chosen parametric values Fig. (2), we clearly observe thermodynamically stable behavior for the model within the investigated regime. This analysis reinforce the black hole nature of the configuration and lays the foundation for studying phenomena such as non-monotonic angular velocity profile in the context of Aschenbach-like effects. We now turn to its geodesic behavior of the model. As previously noted, by examining the general form of the metric function as outlined in Eq. (2.1), one can predict two distinct regimes for characteristic behaviors of TCOs based on the stability of the photon sphere:
1) For radial values $r<r_{\text{PS}}$, the quantity $\beta$ becomes negative, indicating that TCOs are forbidden in this domain. Consequently, the allowed spatial region for TCOs is restricted to radii $r>r_{\text{PS}}$.
2) In contrast, for $r>r_{\text{PS}}$, a negative $\beta$ similarly prohibits the existence of TCOs. Here, TCOs are confined to the interior region with radii $r<r_{\text{PS}}$. This geometric restriction reflects the causal and dynamical influence of the photon sphere on the surrounding geodesic structure. The sign and behavior of $\beta$ act as a discriminant for admissible orbital zones and play a crucial role in shaping the morphology of accretion disks and observational features such as photon rings and ISCO-related dynamic.
Considering the Eq. (2.2) and Eq. (2.5) we obtain the effective potential $H$ and components of the vector field $\phi(r,\Theta)$ as:
$H=\frac{\sqrt{1-\frac{2r^{2}M}{\left(q^{2}+r^{2}\right)^{\frac{3}{2}}}+\frac{r^{2}}{l^{2}}+\left(\frac{1}{2}rc_{1}C+C^{2}c_{2}\right)m_{g}^{2}}}{\sin\!\left(\theta\right)r},$ (3.6)
$\phi^{r_{h}}=-\frac{\left(\left(c_{2}m_{g}^{2}C^{2}+\frac{1}{4}rc_{1}m_{g}^{2}C+1\right)\left(q^{2}+r^{2}\right)^{\frac{5}{2}}-3M\,r^{4}\right)\csc\!\left(\theta\right)}{\left(q^{2}+r^{2}\right)^{\frac{5}{2}}r^{2}},$ (3.7)
$\phi^{\Theta}=-\frac{\sqrt{\frac{\left(\left(c_{2}m_{g}^{2}C^{2}+\frac{1}{2}rc_{1}m_{g}^{2}C+1\right)l^{2}+r^{2}\right)\left(q^{2}+r^{2}\right)^{\frac{3}{2}}-2r^{2}M\,l^{2}}{\left(q^{2}+r^{2}\right)^{\frac{3}{2}}l^{2}}}\,\cot\!\left(\theta\right)\csc\!\left(\theta\right)}{r^{2}}.$ (3.8)
Figure 3: (3a): The photon spheres location at $(r,\theta)=(7.7667,1.57)$ and $(r,\theta)=(12.0996,1.57)$ with respect to $M=2,l=1,q=0.5,C=1,c_{1}=-1,c_{2}=1,m_{g}=0.5$ in the $(r-\theta)$ plane of the normal vector field $n$ , (3b): the topological potential H(r) for Bardeen black holes in massive gravity
As illustrated in Fig. (3), while the model indeed exhibits black hole behavior, a notable feature emerges: two distinct photon spheres —an unstable photon sphere (U-PS:corresponding to a maximum of the potential $H$) and a stable photon sphere (S-PS:associated with a local minimum of HH)—both appear outside the event horizon. Now, if we look at the beta function for this case Fig. (4), we will see that, as mentioned, the region between these two photon spheres has positive beta, and outside it, as stated in the two conditions above, there are forbidden regions due to the negative beta.
Figure 4: (4a): $\beta$ diagram (4b):MSCO localization and space classification for the Bardeen black holes in massive gravity
In Fig. (4), For potential $V$,as defined by Eq. (2.9) we plot the second derivative of the $V$, to investigate the geodesic behavior of spacetime surrounding the black hole. The profile reveals the presence of two marginally stable circular orbits (MSCOs) :
- The first MSCO, located within the physically admissible domain, marks the transition between stable and unstable TCOs. Its presence serves as the boundary for geodesic stability in orbital motion. This structure creates a geodesically bounded region between the two photon spheres where stable circular motion is allowed—a behavior that strongly correlates with the conditions necessary for the emergence of Aschenbach-like profiles in the angular velocity.
- The second MSCO, although mathematically identifiable as a transitional point between two distinct orbital regimes, resides in a region where the function $\beta$ becomes negative. Since negative $\beta$ implies imaginary energy or angular momentum for test particles, this zone lacks physical legitimacy and is classified as a forbidden region .
Having established the theoretical framework, physical prerequisites, and resulting conditions, we now turn our attention to the angular velocity profile and examine its behavior within the scope of this model. Wei in his valuable work [15] and we in our previous work [11], thoroughly explored the mechanism underlying the emergence of this phenomenon in static black hole configurations, highlighting the dependencies of angular velocity on several interconnected quantities:
- Energy and angular momentum,
- The form of the metric function,
- The behavior of the $\beta$ parameter, which governs geodesic admissibility.
[…]
This multi-variable interdependence implies that deviations from monotonicity in the angular velocity are not merely geometric artifacts, but rather stem from a delicate balance between metric structure and dynamical constraints.
Figure 5: Angular velocity VS r with $M=2,l=1,q=0.5,C=1,c_{1}=-1,c_{2}=1,m_{g}=0.5$ for the Bardeen black holes in massive gravity
As illustrated in Fig. (5), the angular velocity exhibits a decreasing trend along unstable TCOs originating from the edge of the unstable photon sphere —a behavior that aligns with theoretical expectations.
However, notably, this monotonic decline does not persist indefinitely. Upon approaching the minimum of the $\beta$ function at $r=9.7006$, and continuing toward the stable photon sphere —which structurally coincides with a local minimum in the effective potential —the angular velocity begins to increase, giving rise to a globally non-monotonic profile. This reversal in slope is precisely the type of behavior we aimed to identify as indicative of Aschenbach-like dynamics.
A particularly intriguing feature visible in the figure is that the angular velocity’s decreasing trend does not coincide with the location of the MSCO . Moreover, even some stable TCOs appear to experience a reduction in angular velocity, suggesting that the inflection point of the profile is governed more by the underlying spacetime geometry—particularly the structure of the effective potential and the behavior of $\beta$—than by traditional orbital stability criteria.
This detection strengthens the argument that in the absence of rotational frame-dragging, the emergence of non-monotonicity in angular velocity can still arise due to spacetime curvature alone, provided the geometry allows for appropriate potential extrema. Consequently, the present model, now confirmed to host both a stable photon sphere and geodesically bounded orbital zones outside the event horizon, is an ideal candidate for detecting Aschenbach-like behavior within a non-rotating framework.
3.2 Non-linear charged AdS black hole in massive gravity
If we use the following Lagrangian instead of the Bardeen Lagrangian in Eq. (3.1),
$\mathcal{L}(F)=Fe^{-\frac{k}{2Q}\left(2Q^{2}F\right)^{\frac{1}{4}}},$
we will arrive at a nonlinear charged AdS model in massive gravity with metric of the spacetime as follows [35]:
$f(r)$ $=$ $1-\frac{2M}{r}e^{-\frac{k}{2r}}+\frac{r^{2}}{l^{2}}+m_{g}^{2}\left(\frac{Cc_{1}r}{2}+C^{2}c_{2}\right),$ (3.10)
where k is a fixed characteristic parameter of the non-linear electrodynamics by which the charge $Q$ and the mass $M$ of the system are related as $Q^{2}=Mk$.
Given that the methodology and analytical framework have been thoroughly outlined in preceding sections, and considering the recurring structure of the study, we will refrain from repeating detailed procedural explanations in the remainder of this work. To maintain clarity and avoid unnecessary fatigue or confusion, from this point onward, we shall focus exclusively on presenting essential findings and critical observations relevant to each parametric scenario. The behavioral differences of the metric function, driven by parametric variations, are clearly observable in Fig. (6). It can be clearly seen that while changing $l$ and $c_{1}$ does not have a particular effect in terms of metric and geometric horizon, increasing k causes the model to move towards extremality.
Figure 6: Metric function Fig (1a): With $l=1,k=0.5,C=1,c_{1}=-1,c_{2}=1,m_{g}=0.53$ with different M , (1b): With $M=5,l=1,k=0.5,C=1,c_{1}=-1,c_{2}=1$ with different $m_{g}$ (1c): With $M=5,l=1,C=1,c_{1}=-1,c_{2}=1,m_{g}=0.5,$ with different k (1d): With $M=5,k=0.5,C=1,c_{1}=-1,c_{2}=1,m_{g}=0.5$ with different $l$ (1e): With $M=5,l=1,k=0.5,C=1,c_{2}=1,m_{g}=0.5$ with different $c_{1}$ for Non-linear charged AdS black hole in massive gravity
In search of the Aschenbach-like phenomenon
To proceed toward detecting Aschenbach-like behavior, we now consider a baseline configuration by assigning the values $M=2,l=1,k=0.5,C=1,c_{1}=-1,c_{2}=1,m_{g}=0.53$ to the relevant parameters. By numerically solving the metric function, we find that the spacetime possesses two distinct horizons:
- A Cauchy horizon at $r=0.0643$,
- An event horizon at $r=1.2381$.
Now with respect to Eq. (2.2) and Eq. (2.5) for the effective potential $H$ and components of the vector field $\phi(r,\Theta)$ we have:
$H=\frac{\sqrt{1-\frac{2{\mathrm{e}}^{-\frac{k}{2r}}M}{r}+\frac{r^{2}}{l^{2}}+\left(\frac{rc_{1}}{2}+Cc_{2}\right)Cm_{g}^{2}}}{\sin\!\left(\theta\right)r},$ (3.11)
$\phi^{r_{h}}=-\frac{\left(\frac{M\left(k-6r\right){\mathrm{e}}^{-\frac{k}{2r}}}{2}+r^{2}\left(c_{2}m_{g}^{2}C^{2}+\frac{1}{4}rc_{1}m_{g}^{2}C+1\right)\right)\csc\!\left(\theta\right)}{r^{4}},$ (3.12)
$\phi^{\Theta}=-\frac{\sqrt{2}\,\sqrt{\frac{-4{\mathrm{e}}^{-\frac{k}{2r}}M\,l^{2}+2r\left(\left(c_{2}m_{g}^{2}C^{2}+\frac{1}{2}rc_{1}m_{g}^{2}C+1\right)l^{2}+r^{2}\right)}{r\,l^{2}}}\,\cot\!\left(\theta\right)\csc\!\left(\theta\right)}{2r^{2}}.$ (3.13)
Figure 7: (7a): The photon spheres location at $(r,\theta)=(8.0160,1.57)$ and $(r,\theta)=(9.8766,1.57)$ with respect to $M=2,l=1,k=0.5,C=1,c_{1}=-1,c_{2}=1,m_{g}=0.53$ in the $(r-\theta)$ plane of the normal vector field $n$ , (7b): the topological potential H(r) for Non-linear charged AdS black hole in massive gravity
As can be seen in Fig. (7), for the selected parameters, a potential minimum and a maximum are still evident outside the event horizon, representing a stable sphere photon (red dot) and an unstable sphere photon (green dot) in the $\phi^{r}$ diagram.
Figure 8: (8a): $\beta$ diagram, UPS(bluedot)=8.0160U-PS(bluedot)=8.0160 and SPS(reddot)=9.8766S-PS(reddot)=9.8766 (8b): MSCO localization and space classification for the Non-linear charged AdS black hole in massive gravity
In Fig. (8), we have plotted the second derivative for the potential $V$ as it is evident. It can still be seen that there are two MSCOs, the first one is in the acceptable physical range and the second one is in the negative $\beta$ region and is classified as a forbidden region.
Figure 9: Angular velocity VS $r$ with $M=2,l=1,k=0.5,C=1,c_{1}=-1,c_{2}=1,m_{g}=0.53$ for the Non-linear charged AdS black hole in massive gravity
Similar to the previous model, as shown in Fig. (9), the angular velocity in this model, initially exhibits a decreasing trend along the time-like unstable circular orbits originating from the edge of the unstable photon sphere - a behavior that is in line with theoretical expectations. However, it is noteworthy that this uniform decrease does not continue indefinitely. Approaching the minimum $\beta$ at $r=8.8997$ and continuing towards the stable photon sphere - which structurally corresponds to a local minimum in the effective potential - the angular velocity starts to increase, leading to a globally non-uniform profile. This slope inversion is the hallmark of the Aschenbach-like phenomenon.
3.3 ModMax-dRGT-like Black hole in massive gravity
If we use the following Lagrangian in Eq. (3.1),we will arrive at a ModMax-dRGT-like Black hole in massive gravity with metric of the spacetime as follows [36]:
$\mathcal{L}(\mathcal{F})=\mathcal{S}\cosh\gamma-\sqrt{\mathcal{S}^{2}+\mathcal{P}^{2}}\sinh\gamma,$ (3.14)
$f(r)=1-\dfrac{M}{r}-\frac{\Lambda}{3}r^{2}+\frac{q^{2}e^{-\gamma}}{r^{2}}+m_{g}^{2}C\left(\dfrac{c_{1}r}{2}+c_{2}C\right),$ (3.15)
where $\mathcal{S}=\frac{\mathcal{F}}{4}$ and $\mathcal{P}=\frac{\widetilde{\mathcal{F}}}{4}$, respectively, are a true scalar, and a pseudoscalar and $\mathcal{F}=F_{\mu\nu}F^{\mu\nu}$ is the Maxwell invariant. Also $\Lambda$, is the cosmological constant, $M$ and $\gamma$ are mass and the dimensionless parameter (which is known as the ModMax’s parameter).As can be seen in Fig. (10), since in this model the $\gamma$ parameter will have a significant effect on the gravitational behavior of the model, in such a way that small values (e.g. $\gamma<0.6686$) will push the model towards extremalization and larger values will lead to black hole behavior, we therefore plot the metric function in two cases.
Figure 10: Metric function Fig (10a): With $M=1.5,\Lambda=-0.5,q=1,C=0.4,c_{1}=-12,c_{2}=25,m_{g}=0.1$ with different $\gamma$ for ModMax-dRGT-like Black hole in massive gravity
1.( $\gamma<0.6686$)
As shown in Fig. (11), depending on the parametric choices, the structure exhibits black hole behavior only when the mass of the graviton component approaches the mass of the model. Otherwise, varying the parameters has practically no effect, leaving the model in a singularity state without a horizon.
Figure 11: Metric function Fig (11a): With $M=1.5,\Lambda=-0.5,\gamma=0.3,q=1,C=0.4,c_{1}=-12,c_{2}=25,$ with different $m_{g}$ , (11b): With $M=1.5,\gamma=0.3,q=1,C=0.4,c_{1}=-12,c_{2}=25,m_{g}=0.5$ for different $\Lambda$ (11c): With $M=1.5,\Lambda=-0.5,\gamma=0.3,q=1,C=0.4,c_{2}=25,m_{g}=0.5$ with different $c_{1}$ for the ModMax-dRGT-like Black hole in massive gravity
2.($\gamma>0.6686$)
In Fig. (12), we see that increasing the strength of the electromagnetic field (through increasing $\gamma$) with the same choice of parameters as before causes the behavior of the model to change completely, and it exhibits the event horizon and takes the form of a black hole. In Fig. (12), we can see that increasing the gravitational mass increases the number of horizons.
Figure 12: Metric function Fig (12a): With $M=1.5,\Lambda=-0.5,\gamma=2,q=1,C=0.4,c_{1}=-12,c_{2}=25,$ with different $m_{g}$ , (12b): With $M=1.5,\gamma=2,q=1,C=0.4,c_{1}=-12,c_{2}=25,m_{g}=0.1$ for different $\Lambda$ (12c): With $M=1.5,\Lambda=-0.5,\gamma=2,q=1,C=0.4,c_{2}=25,m_{g}=0.1$ with different $c_{1}$ for the ModMax-dRGT-like Black hole in massive gravity
In search of the Aschenbach-like phenomenon
To proceed toward detecting Aschenbach-like behavior, we now consider a baseline configuration by assigning the values $M=2,\Lambda=-0.5,\gamma=2,q=1,C=0.4,c_{1}=-12,c_{2}=25,m_{g}=0.5$ to the relevant parameters. By numerically solving the metric function, we find that the spacetime possesses two distinct horizons:
- A Cauchy horizon at $r=0.0728$,
- An event horizon at $r=1.2531$.
Again, with considering the Eq. (2.2) and Eq. (2.5) we have:
$H=\frac{\sqrt{9-\frac{9M}{r}-3r^{2}\Lambda+\frac{9q^{2}{\mathrm{e}}^{-\gamma}}{r^{2}}+9\left(\frac{rc_{1}}{2}+Cc_{2}\right)Cm_{g}^{2}}}{3\sin\!\left(\theta\right)r},$ (3.16)
$\phi^{r_{h}}=-\frac{\csc\!\left(\theta\right)\left(4C^{2}r^{2}c_{2}m_{g}^{2}+C\,r^{3}c_{1}m_{g}^{2}+8q^{2}{\mathrm{e}}^{-\gamma}+4r^{2}-6Mr\right)}{4r^{4}},$ (3.17)
$\phi^{\Theta}=-\frac{\sqrt{6}\,\sqrt{\frac{3C\,r^{3}c_{1}m_{g}^{2}-2r^{4}\Lambda+\left(6C^{2}c_{2}m_{g}^{2}+6\right)r^{2}-6Mr+6q^{2}{\mathrm{e}}^{-\gamma}}{r^{2}}}\,\cot\!\left(\theta\right)\csc\!\left(\theta\right)}{6r^{2}}.$ (3.18)
Figure 13: (13a): The photon spheres location at $(r,\theta)=(2.0913,1.57)$ and $(r,\theta)=(4.4789,1.57)$ with respect to $M=2,\Lambda=-0.5,\gamma=2,q=1,C=0.4,c_{1}=-12,c_{2}=25,m_{g}=0.5$ in the $(r-\theta)$ plane of the normal vector field $n$ , (13b): the topological potential H(r) for the ModMax-dRGT-like Black hole in massive gravity
For the selected parameters, a potential extremums are still evident outside the event horizon, representing a stable sphere photon (red dot) and an unstable sphere photon (green dot) in the $\phi^{r}$ diagram, see Fig. (13).
Figure 14: (14a): $\beta$ diagram (14b):MSCO localization and space classification for the ModMax-dRGT-like Black hole in massive gravity
As it is evident in Fig. (14), for the second derivative of potential $V$ there are two MSCOs. The first one is in the acceptable physical range and the second one is in the negative $\beta$ region and is classified as a forbidden region.
Figure 15: Angular velocity VS r with $M=2,\Lambda=-0.5,\gamma=2,q=1,C=0.4,c_{1}=-12,c_{2}=25,m_{g}=0.5$ for the ModMax-dRGT-like Black hole in massive gravity
The angular velocity in this model Fig. (15) also initially exhibits a decreasing trend along the time-like unstable circular orbits and approaching the minimum $\beta$ at $r=8.8997$ and continuing towards the stable photon sphere the angular velocity starts to increase, leading to a globally non-uniform profile.
4 Conclusions
The discovery of a non-monotonic angular velocity profile near rotating Kerr black holes—an observation fundamentally at odds with Newtonian expectations—led to the identification of a phenomenon now known as the Aschenbach effect. This behavior, rooted in two quintessential relativistic features— frame dragging and intense spacetime curvature —has become a striking indicator of general relativity’s dominance over classical physics in strongly curved geometries. This raises a pivotal question:
Is non-monotonicity in angular velocity an exclusive feature of rotating spacetimes, or can it serve as a more universal marker of relativistic behavior, even in static geometries?
This motivated investigations into non-rotating black hole models to explore whether such behavior could emerge independent of frame dragging [11, 15, 16].
In our previous study [11], we examined the mechanism responsible for this phenomenon across various ultra-compact gravitational configurations, both in the form of black holes and naked singularities. We found a consistent feature shared by all examined models [11, 15, 16]: the presence of a stable gravitational potential minimum—either local or global—which corresponds to a stable photon sphere. In naked singularity models, the absence of an event horizon naturally permits such minima. However, in black hole configurations, the existence of a stable minimum depends on its placement outside the event horizon, adding a geometric constraint.
This detection strengthens the argument that in the absence of rotational frame-dragging, the emergence of non-monotonicity in angular velocity can still arise due to spacetime curvature alone, provided the geometry allows for appropriate potential extrema. However, this condition is rare among conventional black hole solutions.
Following the analysis of the 4D Einstein–Gauss–Bonnet Massive Gravity model, alongside results reported in [16], we posed a new question:
Can the architectural framework of black holes in Massive Gravity theories accommodate the necessary conditions for Aschenbach-like phenomena?
To address this, we investigated three distinct black hole models within the context of Massive Gravity. Our findings showed that, for some range of parametric configurations (though not for all parameters), Aschenbach-like behavior was reproducible in each case. This provides compelling support for the idea that electromagnetic interaction coupled with Massive Gravity dynamics can yield black hole architectures that naturally support non-monotonic angular velocity profiles.
Since the concept of a photon sphere is intrinsically tied to the framework of general relativity and lacks any classical analogue, both the original Aschenbach effect in rotating spacetimes and the Aschenbach-like phenomenon in static configurations may be considered observable signatures of general relativistic dynamics.
Finally, it is crucial to emphasize, however, that despite sharing the trait of angular velocity variability:
- In rotating black holes, the phenomenon arises from the interplay of frame dragging and strong gravity.
- In static black holes, it is governed by the presence of a minimum in the effective potential outside the horizon.
This fundamental difference in origin justifies the terminological distinction: we refer to the static case as exhibiting an ”Aschenbach-Like” effect.